# Dissecting Open Edge Computing Platforms: Ecosystem, Usage, and Security Risks

### v2

@@ IV-B Edge Computing Nodes
The temporal evolution of edge nodes. Leveraging the timestamps that exist in each passive DNS record, we are allowed to profile not only the temporal evolution of edge nodes but also the lifetime of each individual edge node. Figure 4(a) presents how the number of daily active edge nodes evolves across the time period between January 2021 and November 2023, while Figure 4(b) presents the number of new edge nodes that daily emerge for the same period. As we can see that both OECPs emerged in early 2021 with a small scale of edge nodes, and then quickly ramped up their pools of edge nodes. And the largest volume of daily active edge nodes for TipTime is 3.9 millions while it is 0.8 million for OneThingCloud. On the other hand, we can also observe a large volume of newly emerging edge node IPs, as shown in Figure 4(b). For instance, TipTime has over 10k new edge node IPs observed every day between June, 2021 and November, 2023. Considering the scale of edge node IPs is stable since late 2021, the continuous emergence of new edge node IPs suggests a non-negligible churning rate. We also profile the lifetime of an edge node. We observe that most edge nodes of both OECPs have a short lifetime, which is very different from traditional cloud servers. For instance, over 70% TipTime edge nodes have the lifetime shorter than 10 days while it is 76% for OnethingCloud edge nodes. For more details, please refer to Appendix A.
TABLE V: The list of CDN services and content providers.
CDN OECP Content Provider
Wangsu CDN TipTime Toutiao
Bilibili CDN OneThing Cloud Bilibili
Baidu CDN OneThing Cloud Haokan Video, Baidu Cloud
Xunlei CDN OneThing Cloud Xunlei
IV-C Edge-Assisted Content Delivery
@@
CDNs and content providers. As aforementioned, all edge computing tasks observed in our study are content delivery tasks which involve the collaboration between CDN services and the open edge computing platforms. In total, we have observed the tasks of 6 different CDN services, among which, 2 are observed on the TipTime platform and 4 are on the OneThingCloud. Also, through analyzing the traffic flows of these content delivery tasks, we have identified 16 upstream content providers that subscribe to one or more of these 6 CDN services and have their content payloads delivered through edge nodes of the two OECPs. The full list of CDN services is listed in Table V along with the respective upstream content providers. Also, all CDN services except for Xingyu CDN are independent from the two OECPs.
@@
CDN nodes observed in edge traffic. Also, as detailed in §III, most traffic flows of CDN tasks show distinguishable features for deciding whether a remote party is a CDN node or not. Leveraging these distinguishing features, we are allowed to profile the scale and distribution of CDN nodes that have ever communicated with edge nodes under our control. Despite being a lower-bound estimate, these CDN nodes turn out to be large-scale and widely distributed, regardless of the CDN services. In total, we have observed over 125K distinct CDN node IPs which feature a wide distribution in 61 /8 IPv4 network blocks and 120 different ASes. For more details, please refer to Appendix B-A.
CDN nodes observed in passive DNS. As aforementioned, some CDN services assign each CDN node with one or more unique FQDNs which follow unified patterns. For instance, all YunFan CDN nodes (including our TipTime edge nodes) will have a FQDN under the pattern of *yf*.free-lbv6.idouyinvod.com , while it is *.vld.szbdyd.com for CDN nodes of Xingyu. It allows us to extensively profile the scale of CDN nodes by looking up passive DNS. Table VI lists for each CDN service, the FQDN patterns of its CDN nodes, the number of unique FQDNs and the IP addresses recorded in passive DNS between May 2016 and November 2023. As we can see, CDN nodes are of a much larger scale than that observed in our edge traffic. In total, we have observed over 4 billion FQDNs and 34 million distinct IP addresses that used to serve as CDN nodes.
TABLE VI: CDN nodes observed in passive DNS.
CDN 1 Node FQDN Pattern FQDNs IPs
YunFan *yf*.nodeedge.cn 601,996,484 8,205,570
*yf*.free-lbv6.idouyinvod.com 1,613,619,196 20,469,328
*yf*.dl.jcloudimg.com 302,535,129 6,162,593
ndsv1-*.jdcloudstatus.net /*yf*.jdcloudstatus.net 761,048,329 16,149,799
ndsv1-*.cdnnode.cn 257,696,943 6,778,555
ndsv1-*.cachenode.cn 696,675,292 12,792,771
All 4,233,571,373 28,212,313
Xingyu *.vld.szbdyd.com /*.v2l.szbdyd.com 100,492,251 7,383,677
Wangsu dg*.ourdvsss.com 370,137 322,859
All All 4,334,434,710 34,597,596
  • 1
    Since the three CDN services have very few nodes observed before May 2016, we set the time window as May, 2016 to November, 2023, when querying passive DNS.
End clients of content delivery. In addition to the interaction with CDN nodes, our edge nodes have also delivered content to many end clients of different content providers (e.g., video streaming services). Also, multiple traffic signatures for such edge-to-client interaction have also been identified, which allows us to locate with high confidence a lower-bound set of end clients as well as the respective content providers, e.g., video viewers and the respective video streaming platforms. In total, 1,069,962 distinct IPs of end clients were observed to have ever retrieved content files from our edge nodes and these end clients are users of 15 different content providers which include popular video/live streaming services (e.g., Douyin and Kuaishou), major ecommerce platforms (JD Cloud), and well-adopted cloud storage services (e.g., Baidu Cloud). For more details, please refer to Appendix B-B.
@@
Content payloads under delivery. Across both OECPs, edge nodes are used to deliver content of diverse categories, which include not only traditional static web files (Javascript and media files), but also emerging content types (program files and machine learning models). Particularly, edge nodes under our control used to deliver files of machine learning models to end users, e.g., .tflite, .model, .mlmodel, and .weights. Further investigation reveals that they are likely used for face-relevant processing for users of Kuaishou. For more details, please refer to Appendix B-C.
The quality of service of content delivery. It is observed that content recipients of an edge node tend to be located nearby the edge node, e.g., 18% content recipients are within the range of 50KM of an edge node deployed in Shanghai, while an edge node located in Shenzhen has over 55% content recipients distributed within a distance of 300KM. Also, compared with traditional CDNs and closed edge platforms, the delay of edge-assisted content delivery is either comparable or even lower. For instance, edge-assisted content delivery can achieve a lower 95-percentile delay of 83.37ms and 93.75ms for YunFan CDN and Baidu CDN respectively, while it is 103.38ms for Akamai and 221.74ms for Edgio [25]. For more details, please refer to Appendix B-E.
@@ V The Security Risks
The threat model. Before diving into security risks, it is necessary to first highlight the threat model upon which we reason about the feasibility and impact of any security risk. First of all, as edge nodes are recruited from the public along with a low vetting bar, we assume some edge nodes can be byzantine or malicious, i.e., under the full control of an attacker. Then, when an edge node under the attacker’s control serves as a CDN node, we assume the attacker has full access to the content payloads under delivery. Although many content payloads can be either encrypted or proprietorially encoded when getting stored in local disks, we observe that such operations of encoding or encryption are conducted on the edge side and are thus visible to the edge node. However, although an attacker can modify the content payloads that are delivered through the edge node under its control, we argue that such a modification can be easily detected and the modified payloads can be discarded. This is because CDN services and content providers can easily instruct the end-user devices to conduct integrity checking for content payloads received from edge nodes and the materials of integrity checking (e.g., signatures, or messaging authentication code) can be transmitted through other secure channels, e.g., a direct secure connection between the end-user device and the server of the content provider.
V-A The Security Testbed
@@
Due to aforementioned ethical considerations, our MITM experiments have a limited coverage of edge traffic flows, i.e., only the edge-wide control-plane traffic and the CDN-specific control-plane traffic are tested. Also, our MITM experiments consider only the most common but simple scenario wherein the attacker sitting at an intermediate hop utilizes a self-signed root certificate to dynamically issue leaf certificates for the requested domain name. Despite these limitations, we are surprised to observe that edge nodes and CDN modules across the two OECPs fail to validate the TLS certificate for many critical traffic flows, for which, more details will be presented in §V-D.
V-B The Exposure of Credentials to Attackers
@@
One more issue resides in the distribution of these credential materials. Specifically, when the YunFan CDN module is instantiated on an edge node, it will be instructed to download the TLS credentials from the control server. Although the payload is transmitted over HTTPS, the credential hosting server has enforced no authentication for downloading requests and the downloading URLs have also got leaked in the plaintext logging traffic towards the control server, which allows any intermediate hops to learn these downloading URLs from passively observing the plaintext logging traffic and then download the TLS credentials even without direct control of a running edge node. We first observed this issue in late 2022, which had then been fixed by YunFan in early 2023. However, due to the failure of TLS certificate verification as detailed below in §V-D, an attacker can still learn the credential downloading URLs even if it has no direct access to any edge nodes.
For all the issues and impact we list above, we discuss the solutions in §VI, e.g, adopt the short-lived and node-specific credentials.
V-C Low Threat Reputation of Edge Node IPs
@@
TABLE VII: The threat reputation of edge node IPs as revealed by the proprietary threat platform.
@@
We thus move to quantitatively profile the threat reputation of edge node IPs, which is achieved through analyzing threat reports aggregated by two representative threat intelligence platforms. One is VirusTotal [30], a global and publicly available threat intelligence platform. Then, considering most edge nodes are located in China, the other platform under our consideration is a proprietary one maintained by a top security vendor that is dedicated to threat intelligence in China. Then, due to rate limits enforced by both platforms as well as the large scale of observed edge nodes, a random sampling strategy was applied to edge node IPs when querying both platforms. As discussed in §IV, the captured edge node IPs belong to different groups, depending on the respective edge platform, whether they were captured from the raw traffic or the passive DNS, and whether they have been confirmed to be edge nodes or just CDN nodes (i.e., potential edge nodes). For each OECP, we define three groups of node IPs. One is the group of edge node IPs as observed in edge traffic, which we name as $\text{Edge}_{\text{Traffic}}$. Then, it is $\text{CDN}_{\text{Traffic}}$ which encompasses IP addresses of CDN nodes as observed in edge traffic. Lastly, the third group $\text{Edge}_{\text{DNS}}$ consists of edge node IP addresses as learned from passive DNS. Then, to query both threat intelligence platforms, we randomly sampled up to 10K IPs from each of all these groups for each OECP.
As learned from documentations of both threat intelligence platforms, they have different advantages in terms of profiling malicious activities of a given IP address. Specifically, VirusTotal focuses more on how an IP acts as a server in malicious activities, such as hosting malware or phishing websites. In contrast, the private threat platform is very helpful to profile how an IP serves as a malicious client when involving malicious activities, e.g., contacting a botnet C2 server. Therefore, we separately present the threat results as revealed by these two platforms.
@@
Threat reports from the proprietary threat intelligence platform. We first look into malicious traces of edge nodes as learned from the proprietary threat intelligence platform, which reveals that edge node IPs are concurrently involved in malicious activities that feature both a large scale and diverse categories. As shown in Table VII, a large portion of edge node IPs have been associated with malicious traffic flows (MTFs) that were captured by the proprietary threat dataset between January, 2022 and November, 2023. Particularly, among TipTime edge nodes directly observed in edge traffic, 79.66% have been involved in 10 or more MTFS, while it is 45.97% for OneThingCloud edge nodes. Also, among edge node IPs associated with one or more MTFs, the median number of MTFs is 275, while the average number of MTFs is 34,227. Even if we increase the bar for associating an IP with MTFs through counting only MTFs that are captured on the exact date when an IP serves as an edge node, we can still observe such an extensive involvement in malicious activities. While only counting MTFs captured for dates when an IP serves as an edge node, 6.15% TipTime edge nodes are still associated with one or more MTFs while it is 2.94% for OneThingCloud edge nodes. For more details, please refer to Table XX in Appendix C.
TABLE VIII: Top 5 categories of malicious traffic flows.
@@
We then take a closer look into categories of the edge-involved MTFs. In total, the proprietary threat platform has observed 13 different categories of MTFs. Table VIII presents top 5 along with their contribution to MTFs and the involved edge nodes. These top five categories include botnet, remote access trojan (RAT), illicit promotion, cryptojacking, and malicious downloads. Particularly, over 1.3 billion botnet traffic flows have been captured, which involve 11% of all the sampled edge node IPs. On the other hand, 55.90% edge node IPs are involved in MTFs of RAT which suggest that one or more machines attached to these IPs are compromised with RATs installed. For instance, 48% edge node IPs have involved in traffic flows towards the C2 server (i.e., pro.csocools.com) of DoubleGuns, a RAT campaign being active in the last few years [31].
@@
TABLE IX: Statistics of edge nodes’ malicious activities as captured by VirusTotal: Mal denotes malicious.
Edge Group Mal W Mal URLs W Malware
TipTime $\text{Edge}_{\text{Traffic}}$ 3.99% 3.96% 0.81%
$\text{CDN}_{\text{Traffic}}$ 3.62% 3.59% 0.73%
$\text{Edge}_{\text{DNS}}$ 2.75% 2.73% 0.91%
OneThing $\text{Edge}_{\text{Traffic}}$ 2.33% 2.33% 0.22%
$\text{CDN}_{\text{Traffic}}$ 3.01% 2.99% 0.46%
$\text{Edge}_{\text{DNS}}$ 3.07% 3.04% 1.31%
Both $\text{Edge}_{\text{Traffic}}$ 3.47% 3.44% 0.62%
$\text{CDN}_{\text{Traffic}}$ 3.21% 3.19% 0.55%
$\text{Edge}_{\text{DNS}}$ 2.78% 2.76% 0.94%
As shown in Table IX, a non-negligible fraction of edge node IPs are considered as malicious by VirusTotal for either hosting malicious URLs or distributing malware payloads. Particularly, among edge node IPs directly observed in edge traffic ($\text{Edge}_{\text{Traffic}}$), 3.47% are detected as malicious by VirusTotal. We then look into top IPs that are most reported by VirusTotal, and we observe that 49 out of the top 50 IPs used to be bots of the Mozi IoT botnet since they have one or more URLs detected for Mozi payload distribution. For instance, 183.15.89.96, a TipTime edge node observed in traffic, used to distribute Mozi payloads and has 20 URLs detected by VirusTotal, e.g., http://183.15.89.96:38713/Mozi.m, and https://183.15.89.96:38713/mozi.a. Among IPs having malicious URLs detected, over 25% have such kinds of Mozi payload URLs.
In summary, a large portion of edge node IPs are concurrently involved into diverse malicious activities, which can likely render a low threat reputation and thus disrupt the legitimate edge computing tasks. For example, the CDN domain cdn.thunderstore.io was blocked because a malicious user had uploaded a crypto miner to this site [32].
V-D The Validation Failures of TLS Certificates
TABLE X: The issue of certificate validation failures for edge nodes, wherein ✘ denotes validation failures.
@@ V-D The Validation Failures of TLS Certificates
Aforementioned security risks are inherent in the design of the edge platforms, while the failure of validating TLS certificates should be attributed to either implementation errors or bad security practices. In a nutshell, we observe and demonstrate that edge nodes of both OECPs fail to verify the server TLS certificate for part of the TLS traffic flows. Although it should be considered as an implementation error, we are surprised to see the significant extent it has shown up in edge computing traffic flows. Specifically, when connecting to upstream servers via TLS, some edge nodes and CDN modules fail to properly validate the certificates of many TLS servers, which allows a MITM attacker with no knowledge of the server’s private key to break the confidentiality and integrity of the TLS traffic. Such a failure of certificate validation varies across edge platforms and CDN modules as well as the traffic categories (e.g., traffic for logging, and traffic for caching content). As listed in Table X, a TipTime edge node is subject to this vulnerability for all the TLS traffic towards upstream servers, while only the logging traffic of OneThingCloud edge nodes shares this vulnerability. However, one thing to note, our certificate validation test (§V-A) considers only a simple attacker that is equipped with a self-signed root certificate and can thus serve as a lower-bound estimator for the TLS security risks. In another word, even if the edge node had passed our test, it may still be vulnerable to more complicated and more advanced MITM attacks, e.g., a complicated MITM attacker may instead use a certificate issued from a legitimate CA to domains under its control, in which case, the edge node would still be vulnerable to MITM attacks when it fails to verify the field of the common name.
Then, when it comes to CDN modules, as listed in Table XI, the CDN TLS traffic towards upstream servers can be divided into four categories depending on the the purposes, namely, control-plane communication, caching content from the upstream content servers, logging, and updating TLS credentials (e.g., certificate update). Among all the 6 CDN modules, 4 have adopted the standard TLS protocols for communication with the upstream servers, all of which suffer from the certificate invalidation risk for one or more categories of TLS traffic. Particularly, the YunFan CDN is vulnerable to certificate invalidation for all the four categories of TLS traffic. One thing to note, all these certificate invalidation results have been confirmed by our ethical MITM experiments.
TABLE XI: The issue of certificate validation failures for CDN modules, wherein ✘ denotes validation failures.
@@
V-E Privacy Risks Against End Users

### v3

@@ 4.2 Edge Computing Nodes
The temporal evolution of edge nodes. Leveraging the timestamps that exist in each passive DNS record, we are allowed to profile not only the temporal evolution of edge nodes but also the lifetime of each individual edge node. Figure 4(a) presents how the number of daily active edge nodes evolves across the time period between January 2021 and November 2023, while Figure 4(b) presents the number of new edge nodes that daily emerge for the same period. As we can see that both OECPs emerged in early 2021 with a small scale of edge nodes, and then quickly ramped up their pools of edge nodes. And the largest volume of daily active edge nodes for TipTime is 3.9 millions while it is 0.8 million for OneThingCloud. On the other hand, we can also observe a large volume of newly emerging edge node IPs, as shown in Figure 4(b). For instance, TipTime has over 10k new edge node IPs observed every day between June, 2021 and November, 2023. Considering the scale of edge node IPs is stable since late 2021, the continuous emergence of new edge node IPs suggests a non-negligible churning rate. We also profile the lifetime of an edge node. We observe that most edge nodes of both OECPs have a short lifetime, which is very different from traditional cloud servers. For instance, over 70% TipTime edge nodes have the lifetime shorter than 10 days while it is 76% for OnethingCloud edge nodes. For more details, please refer to Appendix B.
@@
4.3 Edge-Assisted Content Delivery
@@
CDNs and content providers. As aforementioned, all edge computing tasks observed in our study are content delivery tasks which involve the collaboration between CDN services and the open edge computing platforms. In total, we have observed the tasks of 6 different CDN services, among which, 2 are observed on the TipTime platform and 4 are on the OneThingCloud. Also, through analyzing the traffic flows of these content delivery tasks, we have identified 16 upstream content providers that subscribe to one or more of these 6 CDN services and have their content payloads delivered through edge nodes of the two OECPs. For the full list of CDN services along with the respective upstream content providers, please refer to Table IX in Appendix C. Also, all CDN services except for Xingyu CDN are independent from the two OECPs.
@@
CDN nodes observed in edge traffic. Also, as detailed in §3, most traffic flows of CDN tasks show distinguishable features for deciding whether a remote party is a CDN node or not. Leveraging these distinguishing features, we are allowed to profile the scale and distribution of CDN nodes that have ever communicated with edge nodes under our control. Despite being a lower-bound estimate, these CDN nodes turn out to be large-scale and widely distributed, regardless of the CDN services. In total, we have observed over 125K distinct CDN node IPs which feature a wide distribution in 61 /8 IPv4 network blocks and 120 different ASes. For more details, please refer to Appendix IX.
@@
Content payloads under delivery. Across both OECPs, edge nodes are used to deliver content of diverse categories, which include not only traditional static web files (Javascript and media files), but also emerging content types (program files and machine learning models). Particularly, edge nodes under our control used to deliver files of machine learning models to end users, e.g., .tflite, .model, .mlmodel, and .weights. Further investigation reveals that they are likely used for face-relevant processing for users of Kuaishou. For more details, please refer to Appendix D.
@@ 5 The Security Risks
The threat model. Before diving into security risks, it is necessary to first highlight the threat model upon which we reason about the feasibility and impact of any security risk. First of all, as edge nodes are recruited from the public along with a low vetting bar, which only exists in the OECP paradigms, we assume some edge nodes can be byzantine or malicious, i.e., under the full control of an attacker. Then, when an edge node under the attacker’s control serves as a CDN node, we assume the attacker has full access to the content payloads under delivery. Although many content payloads can be either encrypted or proprietorially encoded when getting stored in local disks, we observe that such operations of encoding or encryption are conducted on the edge side and are thus visible to the edge node. However, although an attacker can modify the content payloads that are delivered through the edge node under its control, we argue that such a modification can be easily detected and the modified payloads can be discarded. This is because CDN services and content providers can easily instruct the end-user devices to conduct integrity checking for content payloads received from edge nodes and the materials of integrity checking (e.g., signatures, or messaging authentication code) can be transmitted through other secure channels, e.g., a direct secure connection between the end-user device and the server of the content provider.
5.1 The Security Testbed
@@
Due to aforementioned ethical considerations, our MITM experiments have a limited coverage of edge traffic flows, i.e., only the edge-wide control-plane traffic and the CDN-specific control-plane traffic are tested. Also, our MITM experiments consider only the most common but simple scenario wherein the attacker sitting at an intermediate hop utilizes a self-signed root certificate to dynamically issue leaf certificates for the requested domain name. Despite these limitations, we are surprised to observe that edge nodes and CDN modules across the two OECPs fail to validate the TLS certificate for many critical traffic flows, for which, more details will be presented in §5.4.
5.2 The Exposure of Credentials to Attackers
@@
One more issue resides in the distribution of these credential materials. Specifically, when the YunFan CDN module is instantiated on an edge node, it will be instructed to download the TLS credentials from the control server. Although the payload is transmitted over HTTPS, the credential hosting server has enforced no authentication for downloading requests and the downloading URLs have also got leaked in the plaintext logging traffic towards the control server, which allows any intermediate hops to learn these downloading URLs from passively observing the plaintext logging traffic and then download the TLS credentials even without direct control of a running edge node. We first observed this issue in late 2022, which had then been fixed by YunFan in early 2023. However, due to the failure of TLS certificate verification as detailed below in §5.4, an attacker can still learn the credential downloading URLs even if it has no direct access to any edge nodes.
For all the issues and impact above, we discuss the solutions in Appendix F, e.g, adopt the short-lived and node-specific credentials.
5.3 Low Threat Reputation of Edge Node IPs
@@
TABLE V: The threat reputation of edge node IPs as revealed by the proprietary threat platform.
@@
We thus move to quantitatively profile the threat reputation of edge node IPs, which is achieved through analyzing threat reports aggregated by two representative threat intelligence platforms. One is a proprietary threat intelligence platform maintained by a top security vendor, considering most edge nodes are located in China. Then the global and publicly available VirusTotal [28] is also considered. VirusTotal focuses more on how an IP acts as a server in malicious activities, such as hosting malware or phishing websites. While the private threat platform is very helpful to profile how an IP serves as a malicious client when involving malicious activities, e.g., contacting a botnet C2 server. Then, due to rate limits enforced by both platforms as well as the large scale of observed edge nodes, a random sampling strategy was applied to edge node IPs when querying both platforms. As discussed in §4, the captured edge node IPs belong to different groups, depending on the respective edge platform, whether they were captured from the raw traffic or the passive DNS, and whether they have been confirmed to be edge nodes or just CDN nodes (i.e., potential edge nodes). For each OECP, we define three groups of node IPs. One is the group of edge node IPs as observed in edge traffic, which we name as $\text{Edge}_{\text{Traffic}}$. Then, it is $\text{CDN}_{\text{Traffic}}$ which encompasses IP addresses of CDN nodes as observed in edge traffic. Lastly, the third group $\text{Edge}_{\text{DNS}}$ consists of edge node IP addresses as learned from passive DNS. Then, to query both threat intelligence platforms, we randomly sampled up to 10K IPs from each of all these groups for each OECP.
@@
Threat reports from the proprietary threat intelligence platform. We first look into malicious traces of edge nodes as learned from the proprietary threat intelligence platform, which reveals that edge node IPs are concurrently involved in malicious activities that feature both a large scale and diverse categories. As shown in Table V, a large portion of edge node IPs have been associated with malicious traffic flows (MTFs) that were captured by the proprietary threat dataset between January, 2022 and November, 2023. Particularly, among TipTime edge nodes directly observed in edge traffic, 79.66% have been involved in 10 or more MTFS, while it is 45.97% for OneThingCloud edge nodes. Also, among edge node IPs associated with one or more MTFs, the median number of MTFs is 275, while the average one is 34,227. Even if we increase the bar for associating an IP with MTFs through counting only MTFs that are captured on the exact date when an IP serves as an edge node, we can still observe such an extensive involvement in malicious activities. While only counting MTFs captured for dates when an IP serves as an edge node, 6.15% TipTime edge nodes are still associated with one or more MTFs while it is 2.94% for OneThingCloud edge nodes. For more details, please refer to Table XVI in Appendix E.
TABLE VI: Top 5 categories of malicious traffic flows.
@@
We then take a closer look into categories of the edge-involved MTFs. In total, the proprietary threat platform has observed 13 different categories of MTFs. Table VI presents top 5 along with their contribution to MTFs and the involved edge nodes. These top five categories include botnet, remote access trojan (RAT), illicit promotion, cryptojacking, and malicious downloads. Particularly, over 1.3 billion botnet traffic flows have been captured, which involve 11% of all the sampled edge node IPs. On the other hand, 55.90% edge node IPs are involved in MTFs of RAT which suggest that one or more machines attached to these IPs are compromised with RATs installed. For instance, 48% edge node IPs have involved in traffic flows towards the C2 server (i.e., pro.csocools.com) of DoubleGuns, a RAT campaign being active in the last few years [29].
@@
When it comes to VirusTotal, a non-negligible fraction of edge node IPs are considered as malicious for either hosting malicious URLs or distributing malware payloads. For more details, please refer to Appendix E. In summary, a large portion of edge node IPs are concurrently involved into diverse malicious activities, which can likely render a low threat reputation and thus disrupt the legitimate edge computing tasks. For example, the CDN domain cdn.thunderstore.io was blocked because a malicious user had uploaded a crypto miner to this site [30].
5.4 The Validation Failures of TLS Certificates
TABLE VII: The issue of certificate validation failures for edge nodes, wherein ✘ denotes validation failures.
@@ 5.4 The Validation Failures of TLS Certificates
Aforementioned security risks are inherent in the design of the edge platforms, while the failure of validating TLS certificates should be attributed to either implementation errors or bad security practices. In a nutshell, we observe and demonstrate that edge nodes of both OECPs fail to verify the server TLS certificate for part of the TLS traffic flows. Although it should be considered as an implementation error, we are surprised to see the significant extent it has shown up in edge computing traffic flows. Specifically, when connecting to upstream servers via TLS, some edge nodes and CDN modules fail to properly validate the certificates of many TLS servers, which allows a MITM attacker with no knowledge of the server’s private key to break the confidentiality and integrity of the TLS traffic. Such a failure of certificate validation varies across edge platforms and CDN modules as well as the traffic categories (e.g., traffic for logging, and traffic for caching content). As listed in Table VII, a TipTime edge node is subject to this vulnerability for all the TLS traffic towards upstream servers, while only the logging traffic of OneThingCloud edge nodes shares this vulnerability. However, one thing to note, our certificate validation test (§5.1) considers only a simple attacker that is equipped with a self-signed root certificate and can thus serve as a lower-bound estimator for the TLS security risks. In another word, even if the edge node had passed our test, it may still be vulnerable to more complicated and more advanced MITM attacks, e.g., a complicated MITM attacker may instead use a certificate issued from a legitimate CA to domains under its control, in which case, the edge node would still be vulnerable to MITM attacks when it fails to verify the field of the common name.
Then, when it comes to CDN modules, as listed in Table VIII, the CDN TLS traffic towards upstream servers can be divided into four categories depending on the the purposes, namely, control-plane communication, caching content from the upstream content servers, logging, and updating TLS credentials (e.g., certificate update). Among all the 6 CDN modules, 4 have adopted the standard TLS protocols for communication with the upstream servers, all of which suffer from the certificate invalidation risk for one or more categories of TLS traffic. Particularly, the YunFan CDN is vulnerable to certificate invalidation for all the four categories of TLS traffic. One thing to note, all these certificate invalidation results have been confirmed by our ethical MITM experiments.
TABLE VIII: The issue of certificate validation failures for CDN modules, wherein ✘ denotes validation failures.
@@
5.5 Privacy Risks Against End Users